OpenStudy (superhelp101):

How many atoms are in 2.70 moles of iron (Fe) atoms? 4.48 × 10-24 atoms Fe 1.63 × 1024 atoms Fe 2.23 × 1023 atoms Fe 6.02 × 1023 atoms Fe

OpenStudy (superhelp101):

@iPwnBunnies how do you solve this?

OpenStudy (ipwnbunnies):

Use Avogadro's Number. Since we're dealing with atoms of an element. There are 6.02 *10^23 atoms in one mole of an element. Find a conversion factor to find the amount of atoms in 2.70 moles. :3

OpenStudy (superhelp101):

wait how do you know to use Avogadro's number?

OpenStudy (ipwnbunnies):

Just because of the situation. They just ask how many atoms are in 2.70 moles of iron. It doesn't matter which element it is, there will always be 6.02 * 10^23 particles in a mole.

OpenStudy (superhelp101):

ohh. ok let me solve it

OpenStudy (superhelp101):

the second one correct?

OpenStudy (ipwnbunnies):

Yesh. \[2.70 moles * \frac{6.02 * 10^{23} atoms}{mol} = 1.63 * 10^{24} atoms\]

OpenStudy (superhelp101):

yay! thank you!